OpenStudy (anonymous):

Chris bought a package of loose leaf paper that contained 500 sheets. In the first week, he used 1/5 of the paper. During the next two weeks, chris used 2/5 of the remaining paper. How many sheets of paper did chris have available at the start of the fourth week? Please help ))))): i got 40 sheets of paper was my answer but i got it weong so i'm doing corrections. Can you even show me how you did it? I can't respond because the comment bar thing doesn't doesn't have a submit or done button. Sorry!

OpenStudy (anonymous):

started with 500 sheets In the first week he used 1/5....1/5 of 500 = 1/5(500) = 500/5 = 100 Chris used 100 sheets in first week leaving him with 400. He then used 2/5 of remaining paper...2/5 of 400 = 2/5(400) = 800/5 = 160. During those two weeks he used 160 sheets leaving him with 240 sheets. So at the start of the 4th week, he had 240 sheets.

OpenStudy (anonymous):

any questions ?